Two bodies, of mass 50kg and 100kg, are 25cm apart. Calculate the universal gravitational constant G=6.67×10-^11Nm2/kg2 and radius of the earth is 6.4×10^6m.


1
Expert's answer
2021-02-28T07:32:30-0500
"F=\frac{GmM}{r^2}\\F=\frac{(6.67\cdot10^{-11})(50)(100)}{0.25^2}=5.3\cdot10^{-6}N"